# Repeated Padding for Sequential Recommendation

### v2

@@ Abstract.
Sequential recommendation aims to provide users with personalized suggestions based on their historical interactions. When training sequential models, padding is a widely adopted technique for two main reasons: 1) The vast majority of models can only handle fixed-length sequences; 2) Batch-based training needs to ensure that the sequences in each batch have the same length. The special value 0 is usually used as the padding content, which does not contain the actual information and is ignored in the model calculations. This common-sense padding strategy leads us to a problem that has never been explored in the recommendation field: Can we utilize this idle input space by padding other content to improve model performance and training efficiency further?
In this paper, we propose a simple yet effective padding method called Repeated Padding (RepPad). Specifically, we use the original interaction sequences as the padding content and fill it to the padding positions during model training. This operation can be performed a finite number of times or repeated until the input sequences’ length reaches the maximum limit. Our RepPad can be considered as a sequence-level data augmentation strategy. Unlike most existing works, our method contains no trainable parameters or hyperparameters and is a plug-and-play data augmentation operation. Extensive experiments on various categories of sequential models and five real-world datasets demonstrate the effectiveness and efficiency of our approach. The average recommendation performance improvement is up to 60.3% on GRU4Rec and 24.3% on SASRec. We also provide in-depth analysis and explanation of what makes RepPad effective from multiple perspectives. Our datasets and codes are available at https://github.com/KingGugu/RepPad.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: 18th ACM Conference on Recommender Systems, October 14–18, 2024, Bari, Italy; October 14–18,2024; Bari, Italyisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
Sequential recommendation (SR) focuses on characterizing users’ dynamic preferences in their historical sequences to predict the next user-item interaction(s). In the last few years, various sequential models have been proposed to model the transition pattern from the previous item to the next in user sequences. Pioneering works employed Markov Chains (Rendle et al., 2010; Shani et al., 2005) and session-based KNN (He and McAuley, 2016; Hu et al., 2020) to model sequential data. Later, CNN-based (Yuan et al., 2019; Tang and Wang, 2018) and RNN-based (Hidasi et al., 2015a; Liu et al., 2016; Qu et al., 2022) models were proposed to capture the dependencies among items within a sequence or across different sequences. More recently, the powerful capability of Transformers (Vaswani et al., 2017; Devlin et al., 2018) in encoding sequences has been adopted for applications in SR. Transformer-based SR models (Kang and McAuley, 2018; Sun et al., 2019; Li et al., 2020; Dang et al., 2023b, a) encode sequences by learning item importance and behavior relevance via self-attention mechanisms.
@@
Although numerous models with different architectures have been proposed, padding has never been explored as a necessary technique for training sequential models. The vast majority of sequence models can only support inputting fixed-length sequences to define the model architecture and process model calculations. In addition, during training, the Batch Processing techniques require all sequences in each batch to be the same length. Therefore, the maximum sequence length is usually used as a specified hyperparameter before the model training. During training, interceptions are performed for long sequences, and padding is performed for short sequences. Due to user habits and the long-tail effect, only a few users have long interaction sequences, while most users have very short ones (Jiang et al., 2021). When padding a sequence, the special value 0 is usually treated as content. It does not contain meaningful information and is not involved in model calculations. This universally observed convention and custom leads us to the question: Can we utilize this idle input space by padding other content to improve model performance and training efficiency further?
@@
Motivated by the question above, we propose a simple yet effective padding strategy called Repeated Padding (RepPad for short) in this paper. Unlike the traditional padding strategy, we iteratively pad the original whole interaction sequence. This operation can be performed a limited number of times or repeated until the sequence length reaches a preset maximum limit. To avoid the situation where the head of the sequence is used to predict the tail, we add a delimiter between each padding. We aim to utilize these input spaces that are not involved in model training but occupying positions to improve model performance and training efficiency.
@@
From another perspective, RepPad is a sequence-level data augmentation method. The training of recommendation models relies heavily on large-scale labeled data. Data sparsity has been one of the major problems faced by recommender systems (Jing et al., 2023). Earlier works use random data augmentation methods to increase the size of the training dataset directly, such as Dropout (Tan et al., 2016a) and Slide Windows (Tang and Wang, 2018). Besides, some researchers utilize counterfactual thinking (Wang et al., 2021), bidirectional Transformer (Jiang et al., 2021; Liu et al., 2021b), or diffusion model (Liu et al., 2023) to augment sequence data. Inspired by recent developments of Contrastive Learning (CL) (Jing et al., 2023; Yu et al., 2023; Zhao et al., 2023a, b), many effective data augmentation operators have been proposed (Xie et al., 2022; Liu et al., 2021a; Dang et al., 2023b; Qin et al., 2023; Tian et al., 2023). These models typically leverage shared encoders to encode original and augmented sequences. The representations of the original sequences are used for the main recommendation task, and the representations of the augmented sequence are used as the input for contrastive learning (Liu et al., 2021a). These methods also use the special value 0 as the padding content when dealing with sequences whose length is less than a specified value. However, adding additional sequence data to the original dataset makes the training process more time-consuming. Our RepPad utilizes the idle padding space to perform repeated padding before the sequences are fed into the model. There is no additional increase in the number of sequences that need to be involved in training.
@@
In the discussion section, we analyze the advantages of RepPad over existing methods and present possible drawbacks. Different from elaborate data augmentation operators or the methods that require training, our method requires no training process and introduces no additional hyper-parameters or trainable parameters. It is independent of the backbone network and is a data augmentation plugin. We implement RepPad on various representative models based on different architectures. Extensive experimental results demonstrate the effectiveness and efficiency of our method. It also achieves satisfactory improvements over existing data augmentation methods. Furthermore, we explore the performance of different variants of RepPad. To explore what makes RepPad effective, we provide in-depth analyses from the perspectives of loss convergence and gradient stability.
In summary, our work makes the following contributions:
  • We propose a simple yet effective padding paradigm called Repeated Padding. It does not contain any parameters and is a plug-and-play data augmentation method.
@@
  • Comprehensive experiments on multiple datasets, various types of sequential models, and augmentation methods demonstrate the effectiveness and efficiency of RepPad.
@@
  • We provide analysis and insights of why RepPad is effective from the perspectives of loss convergence and gradient stability. To our knowledge, this is the first work to explore the padding strategy in the recommendation field.
@@ 2. Methodology
In this section, we formally present the problem of sequential recommendation. After that, we introduce the current widely used padding strategy and briefly describe the sequence data augmentation. More data augmentation methods will be introduced in section 4, Related Work. Finally, we present our proposed RepPad.
@@ 2.1. Problem Formulation
Suppose we have user and item sets denoted by symbols $\mathcal{U}$ and $\mathcal{V}$, respectively. Each user $u\in\mathcal{U}$ is associated with a sequence of interacted items in chronological order $s_{u}=[v_{1},\ldots,v_{j},\ldots,v_{\left|s_{u}\right|}]$, where $v_{j}\in\mathcal{V}$ indicate the item that user $u$ has interacted with at time step $j$ and $\left|s_{u}\right|$ is the sequence length. The sequential recommendation task can be formulated as follows. Given the sequences of interacted items $s_{u}$, our task is to accurately predict the most possible item $v^{*}$ that user $u$ will interact with at time step $\left|s_{u}\right|+1$, formulated as follows:
@@
This equation can be interpreted as calculating the probability of all candidate items and selecting the highest one for recommendation.
Figure 1. An illustration of (a) the original batch sequences, (b) the traditional padding, (c) traditional padding with data augmentation, and (d-e) our proposed RepPad. We set the batch size = 4 and the maximum sequence length = 10 for demonstration. The $\left|\mathcal{U}\right|$ is the number of users in the dataset.
@@ 2.2. Padding in Sequential Recommendation
Padding is a common technique when training sequential models. The Batch Processing requires that the sequences in each batch have the same length. In addition, common CNN- or Transformer-based models can only handle fixed-length sequences and require model initialization and operations based on fixed sequence lengths. So, before training a recommendation model, we usually specify a maximum sequence length $N$. For sequences with an original length of more than $N$, the most recent $N$ items are usually chosen because they represent the user’s recent preferences. As shown in Figure1 (a) and (b), original sequences of length less than $N$ are padded with the special value 0 to reach the length $N$. The special value 0 contains no actual information and does not participate in model calculations. The above process can be formulated as follows:
@@ 2.3. Sequence Data Augmentation
To alleviate the data sparsity problem, data augmentation is often used to augment the original data to increase the number of sequence involved in model training. Given an original sequence $s_{u}$, a data augmentation operation makes minor but appropriate changes to $s_{u}$ to generate a new sequence $s_{u}^{d}$. Both $s_{u}$ and $s_{u}^{d}$ are used for model training or self-supervised learning. For example, as shown in Figure1 (c), for an original sequence $s_{u}=[v_{5},v_{2},v_{6},v_{4},v_{1},v_{7},v_{9},v_{2}]$, we intercept a portion of the consecutive sequence to generate a new sequence data $s_{u}^{d}=[v_{6},v_{4},v_{1},v_{7},v_{9}]$. For original sequence $s_{u}=[v_{3},v_{7},v_{9}]$, we inert an item $v_{4}$ between $v_{7}$ and $v_{9}$ to generate sequence data $s_{u}^{d}=[v_{3},v_{7},v_{4},v_{9}]$. In each epoch, each sequence will be augmented at least once, so the total number of sequences involved in model training will be at least 2$\left|\mathcal{U}\right|$. The original and augmented sequences also need to apply function Eq. (2) before being fed into the model.
2.4. Ours: Repeated Padding
As illustrated in Figure 1 (b) and (c), a large amount of free space is filled with 0 regardless of whether data augmentation is performed. Also, due to the long-tail effect, the interaction sequences of the vast majority of users are usually short. Therefore, our idea is to utilize these free spaces further to improve the efficiency of model training and recommendation performance.

### v3

# Repeated Padding+: Simple yet Effective Data Augmentation Plugin for Sequential Recommendation

@@ Abstract.
Sequential recommendation aims to provide users with personalized suggestions based on their historical interactions. When training sequential models, padding is a widely adopted technique for two main reasons: 1) The vast majority of models can only handle fixed-length sequences; 2) Batching-based training needs to ensure that the sequences in each batch have the same length. The special value 0 is usually used as the padding content, which does not contain the actual information and is ignored in the model calculations. This common-sense padding strategy leads us to a problem that has never been explored before: Can we fully utilize this idle input space by padding other content to further improve model performance and training efficiency?
In this work, we propose a simple yet effective padding method called Repeated Padding+ (RepPad+). Specifically, we use the original interaction sequences as the padding content and fill it to the padding positions during model training. This operation can be performed a finite number of times or repeated until the input sequences’ length reaches the maximum limit. For those sequences that can not pad full original data, we draw inspiration from the Sliding Windows strategy and intercept consecutive subsequences to fill in the idle space. Our RepPad+ can be viewed as a sequence-level data augmentation strategy. Unlike most existing works, our method contains no trainable parameters or hyperparameters and is a plug-and-play data augmentation operation. Extensive experiments on various categories of sequential models and seven real-world datasets demonstrate the effectiveness and efficiency of our approach. The average recommendation performance improvement is up to 84.11% on GRU4Rec and 35.34% on SASRec. We also provide in-depth analysis and explanation of what makes RepPad+ effective from multiple perspectives.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
As an essential branch of recommender systems, sequential recommendation (SR) has received much attention due to its well-consistency with real-world recommendation situations. It focuses on characterizing users’ dynamic preferences in their historical sequences to predict the next user-item interaction(s). In the last few years, various sequential models have been proposed. Pioneering works employ Markov Chains (Rendle et al., 2010; Shani et al., 2005) and session-based KNN (He and McAuley, 2016; Hu et al., 2020) to model sequential data. Later, CNN (Yuan et al., 2019; Tang and Wang, 2018) and RNN-based (Hidasi et al., 2015; Liu et al., 2016; Qu et al., 2022) models are proposed to capture the dependencies among items within a sequence or across different sequences. More recently, the capability of Transformers (Vaswani et al., 2017; Devlin et al., 2018) in encoding sequences has been adopted for applications in SR. Transformer-based SR models (Kang and McAuley, 2018; Sun et al., 2019; Li et al., 2020; Dang et al., 2023b, a) encode sequences by learning item importance and behavior relevance via self-attention mechanisms.
@@
Although numerous models with different architectures have been proposed, padding has never been explored as an indispensable technique for training sequential models. The vast majority of sequence models can only support inputting fixed-length sequences to define the model architecture and process model calculations. In addition, batch processing techniques require all sequences in each batch to be the same length during training. Therefore, the maximum sequence length $N$ is usually used as a specified hyper-parameter before the model training. During training, interceptions are performed for long sequences with lengths larger than $N$, and padding is performed for short sequences with lengths smaller than $N$. In real-world online platforms, only a few users have long interaction sequences, while most users have very short ones. This phenomenon is also known as the long-tail effect. When padding a sequence, the special value 0 is usually treated as content. It does not contain meaningful information and is not involved in model calculations. As illustrated in Figure 1 (a), this universally observed convention and custom leads us to the question: Can we fully utilize this idle input space by padding other content to improve model performance and training efficiency further?
@@
Motivated by the question above, we propose a simple yet effective padding strategy called Repeated Padding+ (RepPad+ for short), which is based on the previous work RepPad proposed by Dang et al. (Dang et al., 2024). Unlike the traditional padding strategy using special value 0, we iteratively pad the original whole interaction sequence. This operation can be performed a limited number of times or repeated until the sequence length reaches a preset maximum limit. However, some sequences of moderate lengths can make it impossible to fill the remaining space with the complete original sequence. For example, when the original sequence length is 30 and $N=50$, the remaining space of length 20 cannot pad the original sequence. Inspired by the Sliding Windows strategy (Tang and Wang, 2018), we randomly intercept consecutive subsequences from the original sequence to fill these idle spaces. Besides, to avoid the situation where the head of the sequence is used to predict the tail, we add a delimiter between each padding. Our core idea is to utilize these idle input spaces that are not involved in model training but occupy positions to improve model performance and training efficiency.
Figure 1. An illustration of (a) the idle input space in the sequential recommendation and (b) a comparison between traditional SR with padding and our repeated padding.
@@
From another perspective, RepPad+ is a sequence-level data augmentation method. The training of recommendation models relies heavily on large-scale labeled data. Data sparsity has been one of the major problems faced by recommender systems (Jing et al., 2023). Earlier works use random data augmentation methods to increase the size of the training dataset directly, such as Dropout (Tan et al., 2016) and Slide Windows (Tang and Wang, 2018). Besides, some researchers utilize counterfactual thinking (Wang et al., 2021), bidirectional Transformer (Jiang et al., 2021; Liu et al., 2021b), or diffusion model (Liu et al., 2023) to augment sequential data. Inspired by recent developments of Contrastive Learning (CL) (Jing et al., 2023; Yu et al., 2023; Zhao et al., 2023a, b), many effective data augmentation operators have been proposed (Xie et al., 2022; Liu et al., 2021a; Dang et al., 2023b; Qin et al., 2023; Tian et al., 2023). These models typically leverage shared encoders to encode original and augmented sequences. The representations of the original sequences are used for the main recommendation task, and the representations of the augmented sequence are used as the input for contrastive learning. It optimizes encoders by maximizing the agreements between ‘positive’ pairs, which are two augmentations from one sequence (Liu et al., 2021a). As illustrated in Figure 1 (b), these methods also use the special value 0 as the padding content when dealing with sequences whose length is less than a specified value. Besides, adding additional sequence data to the original dataset makes the training process more time-consuming. Our RepPad+ utilizes the idle padding space to perform repetitive padding before the sequences are fed into the model. There is no additional increase in the number of sequences that need to be involved in the training.
@@
Different from elaborate data augmentation operators or the methods that require training, our method requires no training process and introduces no additional hyper-parameters or trainable parameters. It is independent of the backbone network and is a data augmentation plugin. We implement RepPad+ on various representative models based on different architectures. Extensive experimental results demonstrate that our approach can significantly improve the performance of sequential recommendation models based on various types of architectures. Regarding training efficiency and recommendation accuracy, RepPad+ achieves satisfactory improvements over existing heuristic and training-required data augmentation methods. Furthermore, we carefully analyze the different variants of RepPad+. To explore what makes RepPad+ effective, we provide an in-depth analysis from loss convergence, gradient stability, different padding strategies, and future information leakage perspectives.
Overall, our work makes the following contributions:
  • We leverage the idle input space of sequential models and propose a simple yet effective padding paradigm, RepPad+. Compared to RepPad, RepPad+ further considers and augments medium-length sequences.
  • We compared RepPad+ in detail with existing methods. RepPad+ does not contain any parameters, requires no training process, does not increase the data amount in training, and is a plug-and-play augmentation method.
@@
  • We conduct comprehensive experiments on multiple real-world datasets, various architectures of sequential models, and augmentation methods. The results demonstrate the effectiveness and efficiency of our RepPad+.
@@
  • We provide analysis and insights into why RepPad+ is effective from multiple perspectives. To our knowledge, this is the first work to explore the padding strategy in the recommendation system field.
Outline. The rest of this paper is organized as follows. Section 2 provides a brief overview of sequential recommendation models, padding methods in deep learning, and data augmentation for sequential recommendation in the literature. The previous work proposed by Dang et al. (Dang et al., 2024), and our proposed RepPad+ are introduced in Section 3. After that, the experimental results are presented and analyzed in Section 4, followed by the multi-perspective analysis of why our method works in Section 5. Finally, Section 6 concludes our work and outlines future research.
2. Related Works
2.1. Sequential Recommendation
Sequential recommendation (SR) is an emerging topic in the field of recommender systems (Qu et al., 2022). An early solution is to treat the item sequence as a Markov Chain (Rendle et al., 2010; Shani et al., 2005; He and McAuley, 2016), where the next item to predict is closely related to the latest few interactions. The limitation lies in the inability to learn the dependency in a relatively large time step. Hence, a better solution based on recurrent neural networks (RNNs) (Schuster and Paliwal, 1997) emerges since it can capture long- and short-term preferences. A pioneering work GRU4Rec (Hidasi et al., 2015) introduced RNN into the field of recommender systems. It trained a Gated Recurrent Unit (GRU) architecture to model the evolution of user interests. Following this idea, CmnRec (Qu et al., 2022) proposed to divide proximal information units into chunks, whereby the number of memory operations can be significantly reduced. However, RNN cannot be well paralleled and thus will reach its performance bottleneck when dealing with a large volume of sequential data. Another line of research is based on convolutional neural networks (CNNs). The representative work, Caser (Tang and Wang, 2018), embeds a sequence of recent items in the time and latent spaces, then learns sequential patterns using convolutional filters. NextItNet (Yuan et al., 2019) combines masked filters with 1D dilated convolutions to model the dependencies in interaction sequences. More recently, transformer-based approaches have attracted much attention because of their convenience for parallelization and carefully designed self-attention architecture (Vaswani et al., 2017; Devlin et al., 2018). SASRec (Kang and McAuley, 2018) employed unidirectional Transformers to fulfill the next-item prediction task in the sequential recommendation. Some works further optimized the Transformer’s architecture to make it more suitable for recommendation scenarios, such as reducing computational complexity (Fan et al., 2021) and stochastic self-attention (Fan et al., 2022). Furthermore, FMLP-Rec (Zhou et al., 2022) proposes an all-MLP architecture with learnable filters to enhance recommendation performance. Without exception, all of these methods use a zero-padding strategy when training the model. We argue that this strategy yields a huge amount of idle input space, which can be further utilized to improve model performance and training efficiency.
2.2. Padding in Deep Learning
In deep learning, ”padding” refers to the process of adding extra pixels or values around the edges of input data, such as an image or a feature map. It is often used to handle boundary conditions and maintain the spatial dimensions or ensure consistent processing. Padding helps to preserve the integrity and context of the data during convolutional operations and other computations, and it can have a significant impact on the performance and output of deep learning models. For example, zero-padding is a common type where zeros are added. Currently, only a few works have explored the impact of padding on the deep learning models (Islam et al., 2021). An early work (Dwarampudi and Reddy, 2019) analyzed the effects of different padding strategies on Long Short-Term Memory Networks and Convolutional Neural Networks. Further, some researchers demonstrated how the padding mechanism can induce spatial bias in CNNs in the form of skewed kernels and feature-map artifacts (Alsallakh et al., 2020). In addition, some works proposed different padding strategies (Ning et al., 2023), such as leverage partial convolution-based padding (Liu et al., 2018) to improve model performance and training efficiency, distribution padding (Nguyen et al., 2019) to maintain the spatial distribution of the input border data, symmetric padding (Wu et al., 2019) to achieve symmetry within a single convolution layer. The above works focused on the field of computer vision. In addition, the impact of padding in the field of natural language processing has also attracted attention (Gimenez et al., 2020; Nam and Hung, 2019; Cheng, 2020), such as handwriting recognition (Nam and Hung, 2019) and text classification (Cheng, 2020). To our knowledge, we are the first work to explore padding strategy in the sequential recommendation and propose a simple yet effective method.
2.3. Data Augmentation for Sequential Recommendation
Data augmentation has been established as a standard technique to train more robust models in computer vision (Yang et al., 2022; Shorten and Khoshgoftaar, 2019) and natural language processing (Feng et al., 2021; Li et al., 2022). In recommender systems, it has also been widely used and proved to be an effective method. Slide Windows is an early work (Tang and Wang, 2018), which modified the original click sequence by splitting it into many sub-sequences or randomly discarding a certain number of items. Similar to this simple operation, the noise/redundancy injection (Song and Suh, 2022) and subset selection (Tan et al., 2016) have also been proven to be effective enhancement methods. Later, Conditional Generative Adversarial Nets were used to obtain reliable user interaction data from rich side information and further form augmented ground truth training datasets (Wang et al., 2019). Inspired by counterfactual thinking, CASR (Wang et al., 2021) revised the sequence of user behaviors by substituting some previously purchased items with other unknown items. Some researchers proposed to leverage bi-directional transformers (Liu et al., 2021b; Jiang et al., 2021). It employed a reversely pre-trained transformer to generate pseudo-prior items for short sequences. Then, fine-tune the pre-trained transformer to predict the next item. Besides, L2Aug (Wang et al., 2022) categorizes users into core and casual users, using historical data from the former to augment the interaction sequence of the latter. DiffuASR (Liu et al., 2023) leveraged the diffusion model for sequence generation. Two guide strategies are designed to make the generated items correspond to the raw data. With the rise of large language models (LLM), some researchers explored utilizing them as Data Augmenters to improve recommendations for cold-start items (Wang et al., 2024).
In recent years, self-supervised learning techniques (e.g., contrastive learning) (He et al., 2020; Chen et al., 2020) have shown great potential for alleviating the data sparsity problem. Many methods used it as an auxiliary task to further improve sequential model performance (Qin et al., 2023; Xie et al., 2022). For example, CL4SRec (Xie et al., 2022) came up with three data augmentation operators (Crop, Reorder, Mask) for pre-training recommendation tasks equipped with a contrastive learning framework. Based on CL4SRec, CoSeRec (Liu et al., 2021a) integrated item similarity information for data augmentation (Insert, Substitute) with a contrastive learning objective, aiming to maximize the agreement of augmented sequences. To further improve the quality of the augmented sequence, some works performed augmentation concerning time and item category information (Dang et al., 2023a; Tian et al., 2023). In addition to sequence-level augmentation, augments and contrasts at the attribute level (Zhou et al., 2020), embedding level (Qiu et al., 2022), and model level (Hao et al., 2023) has also been explored. Unlike existing methods, we are the first to explore the use of idle padding space for data augmentation. In terms of simplicity and applicability, our method does not contain any parameters or training processes and can be plugged into most existing sequential models.
@@ 3. Methodology
In this section, we formally present the problem of sequential recommendation (Section 3.1). After that, we introduce the current widely used padding strategy (Section 3.2) and briefly describe the sequence data augmentation (Section 3.3). Next, we will present the previous work RepPad, which is proposed by Dang et al. (Dang et al., 2024) (Section 3.4). Based on RepPad, we improve it by considering the medium-length sequences and further propose RepPad+ (Section 3.5). Finally, we provide a discussion between existing and our methods (Section 3.6).
@@ 3.1. Problem Formulation
Suppose we have user and item sets denoted by symbols $\mathcal{U}$ and $\mathcal{V}$, respectively. Each user $u\in\mathcal{U}$ is associated with a sequence of interacted items in chronological order $s_{u}=[v_{1},\ldots,v_{j},\ldots,v_{\left|s_{u}\right|}]$, where $v_{j}\in\mathcal{V}$ indicate the item that user $u$ has interacted with at time step $j$ and $\left|s_{u}\right|$ is the sequence length. The sequential recommendation task can be formulated as follows. Given the sequences items $s_{u}$, how to accurately predict the most possible item $v^{*}$ that user $u$ will interact with at time step $\left|s_{u}\right|+1$, given by:
@@
This equation can be interpreted as calculating the probability of all candidate items and selecting the highest one for the final recommendation.
@@ 3.2. Padding in Sequential Recommendation
Padding is a common technique when training sequential models. The Batch process requires that the sequences in each batch have the same length. In addition, common CNN- or Transformer-based models can only handle fixed-length sequences and require model initialization and operations based on fixed sequence lengths. So, before training a recommendation model, we usually specify a maximum sequence length $N$. For sequences with an original length of more than $N$, the most recent $N$ items are usually chosen because they represent the user’s recent preferences. As shown in Figure 2 (a) and (b), original sequences of length less than $N$ are padded with the special value 0 to reach the length $N$. The special value 0 contains no actual information and does not participate in model calculations. The above process can be formulated as follows:
@@
Figure 2. An illustration of (a) the original batch sequences, (b) the traditional padding, (c) traditional padding with data augmentation, (d,e) the previous work, RepPad, proposed by Dang et al. (Dang et al., 2024), and (f) our proposed RepPad+. We set the batch size = 4 and maximum sequence length = 10 for demonstration. The $\left|\mathcal{U}\right|$ is the number of users in the dataset.
@@ 3.3. Sequence Data Augmentation
To alleviate the data sparsity problem, data augmentation is often used to augment the original data to increase the number of samples involved in model training. Given an original sequence $s_{u}$, a data augmentation operation makes minor but appropriate changes to $s_{u}$ to generate a new sequence $s_{u}^{d}$. Both $s_{u}$ and $s_{u}^{d}$ are used for model training or self-supervised learning. For example, as shown in Figure 2 (c), for an original sequence $s_{u}=[v_{5},v_{2},v_{6},v_{4},v_{1},v_{7},v_{9},v_{2}]$, we intercept a portion of the consecutive sequence to generate a new sequence data $s_{u}^{d}=[v_{6},v_{4},v_{1},v_{7},v_{9}]$. For the original sequence $s_{u}=[v_{3},v_{7},v_{9}]$, we inert an item $v_{4}$ between $v_{7}$ and $v_{9}$ to generate sequence data $s_{u}^{d}=[v_{3},v_{7},v_{4},v_{9}]$.
As previously described, this enhancement process can be accomplished by data augmentation operators, heuristics methods, and training-required synthetic modules. The generated data can be used directly to train recommendation models or for auxiliary tasks. In each epoch, each sequence will be augmented at least once, so the total number of sequences involved in model training will be at least 2$\left|\mathcal{U}\right|$. It is important to note that the original and augmented sequences also need to apply function Eq. (2) before being fed into the model.
3.4. The Previous Work: Repeated Padding (RepPad)
As illustrated in Figure 2 (b) and (c), a large amount of idle space is filled with 0 regardless of whether data augmentation is performed. Also, due to the long-tail effect, the interaction sequences of the vast majority of users are usually short. Therefore, the idea of RepPad is to fully utilize these idle spaces further to improve the efficiency of model training and recommendation performance.
In order to achieve this objective, previous work by Dang et al. (Dang et al., 2024) proposes a simple but effective padding strategy called Repeated Padding (RepPad). Specifically, they utilize the original sequence as padding content instead of the special value 0. As shown in Figure 2 (d), give the original sequence $s_{u}=[v_{1},v_{2},v_{3},v_{4}]$, and they pad the original sequence once on the left and get the padded sequence $s_{u}^{p}=[v_{1},v_{2},v_{3},v_{4},v_{1},v_{2},v_{3},v_{4}]$. This operation can be formulated as:
where $m$ is padding times and |||| indicates sequence concatenation. This operation can be performed only once or repeated several times until there is not enough space left to hold the original sequence. In other words, $m$ can be pre-specified or calculated based on the original sequence length $s_{u}$ and the maximum sequence length $N$. For example, when the maximum sequence length is 10, the remaining length of 2 is insufficient to accommodate the original sequence of length 4, so it still needs to apply function Eq. (2). The final input sequence $s_{u}^{f}=[0,0,v_{1},v_{2},v_{3},v_{4},v_{1},v_{2},v_{3},v_{4}]$. In addition, given the original sequence $s_{u}=[v_{3},v_{7},v_{9}]$, the padding operation can be performed up to 3 times, and the final sequence $s_{u}^{f}=[0,0,v_{3},v_{7},v_{9},v_{3},v_{7},v_{9},v_{3},v_{7},v_{9}]$.
However, there is a problem with the above padding strategy: there may be cases where the end of the original sequence is used to predict the head. Using $[0,0,\underline{v_{1},v_{2},v_{3},v_{4}},\underline{v_{1}},v_{2},v_{3},v_{4}]$ as training data, there will be cases where $v_{1},v_{2},v_{3},v_{4}$ is used to predict $v_{1}$. To solve this problem, they add the special value 0 between each repeated padding to prevent this from happening. This operation can be formulated as:
[…]
As shown in Figure 2 (e), given original sequence $s_{u}=[v_{1},v_{2},v_{3},v_{4}]$, the final sequence after Eq. (4) and (2) is $[0,v_{1},v_{2},v_{3},v_{4},0,v_{1},v_{2},v_{3},v_{4}]$. The possible values of $m$ and their interpretation are given in Table 1.